# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v3

@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{0}}=1$ m is ${\beta_{0}}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\min}}=1$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and ${R_{c}}=600$ m, respectively. The transmit power of each UE is $P_{k}=20$ dBm. Without otherwise stated, the movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-50$ m, ${x_{\max}}={y_{\max}}=50$ m, ${z_{\max}}=40$ m, and $H=500$ m.
@@ V-A Single-Antenna UAV
For further comparison, Fig. 6 shows the minimum achievable rate versus the transmit power of each UE for $K=3$ UEs. Since MMSE beamforming achieves the balance between reducing the interference and noise enhancement, the benchmark scheme of circular array geometry with MMSE beamforming is considered in the following. Moreover, the benchmark scheme of distributed MIMO is considered [1], where four single-antenna access points (APs) cooperatively serve the UEs, and their locations are given by ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, respectively, with ${R_{c}}=1000$ m. It is firstly observed that the minimum achievable rate of both the proposed and benchmark schemes increase as the transmit power increases, as expected. Besides, similar to the case of two UEs, the proposed UAV swarm enabled MA system yields a comparable performance to the IUI-free upper bound, and is significantly superior to the benchmark scheme of circular array geometry with MMSE beamforming. This is expected since the proposed scheme can strike an optimal balance between the beamforming gain improvement and IUI mitigation via the flexible UAV placement position adjustment. It is also observed that the proposed scheme outperforms the benchmark scheme of distributed MIMO as the transmit power increases. This is expected since although the distributed MIMO deploys the geographically separated APs, their fixed-position antennas cannot exploit the channel variation in the spatial domain for mitigating the IUI.
@@ V-B Multi-Antenna UAV
Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the transmit power of each UE is ${P_{k}}=35$ dBm, the UE directions are the same as for Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (21) and (22); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (58); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (21) and (22) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is bounded by the IUI-free communication from above.
@@
Last, Fig. 10 compares the minimum achievable rate for the cases with and without synchronization and position errors, where for the former, the proposed channel estimation-based error compensation scheme is considered. The variances of synchronization and position errors are $1$ ns and ${\sigma_{p}}=\lambda$, respectively. The length of pilot sequence is ${T_{p}}=60$, and the training power is ${p_{{\rm{tr}}}}=P$. It is observed that the proposed channel estimation-based error compensation scheme yields a comparable performance to the case without errors, which demonstrates the effectiveness of the proposed scheme in compensating for synchronization and position uncertainties.
@@ Appendix B Proof of Lemma 1
$={\left({\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}}\right)^{2}}\left[{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{x_{l}^{2}+y_{l}^{2}+z_{l}^{2}-{{\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right)}^{2}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
Let ${C_{\max}}\triangleq\mathop{\max}\limits_{{{\bf{q}}_{l}}\in{\cal C}}x_{l}^{2}+y_{l}^{2}+z_{l}^{2}$, it follows that
@@
$\ \ \ \ \ \left[{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
With ${\nabla^{2}}{g_{k,i,l}}\preceq{\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{2}}{\bf{I}}$, and by choosing ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, we have ${\nabla^{2}}{g_{k,i,l}}\preceq{\delta_{k,i,l}}{\bf{I}}$. Thus, the lower and upper bounds of $g_{k,i,l}$}} in (32) and (33) can be obtained with the Taylor’s theorem [38]. The proof of Lemma 1 is thus completed.

### v4

@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{0}}=1$ m is ${\beta_{0}}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\min}}=1$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and ${R_{c}}=600$ m, respectively. The transmit power of each UE is $P_{k}=20$ dBm. Unless otherwise stated, the movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-50$ m, ${x_{\max}}={y_{\max}}=50$ m, ${z_{\max}}=40$ m, and $H=500$ m.
@@ V-A Single-Antenna UAV
For further comparison, Fig. 6 shows the minimum achievable rate versus the transmit power of each UE for $K=3$ UEs. Since MMSE beamforming achieves the balance between reducing the interference and noise enhancement, the benchmark scheme of circular array geometry with MMSE beamforming is considered in the following. Moreover, the benchmark scheme of distributed MIMO is considered [1], where four single-antenna access points (APs) cooperatively serve the UEs, and their locations are given by ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, ${R_{c}}/2,-{R_{c}}/2,0]^{T}$}, respectively, with ${R_{c}}=1000$ m. It is firstly observed that the minimum achievable rate of both the proposed and benchmark schemes increase as the transmit power increases, as expected. Besides, similar to the case of two UEs, the proposed UAV swarm enabled MA system yields a comparable performance to the IUI-free upper bound, and is significantly superior to the benchmark scheme of circular array geometry with MMSE beamforming. This is expected since the proposed scheme can strike a good balance between the beamforming gain improvement and IUI mitigation via the flexible UAV placement position adjustment. It is also observed that the proposed scheme outperforms the benchmark scheme of distributed MIMO as the transmit power increases. This is expected since although the distributed MIMO deploys the geographically separated APs, their fixed-position antennas cannot exploit the channel variation in the spatial domain for mitigating the IUI.
@@ V-B Multi-Antenna UAV
Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the transmit power of each UE is ${P_{k}}=35$ dBm, the UE directions remain consistent with those in Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (21) and (22); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (58); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (21) and (22) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is upper-bounded by the performance under IUI-free communication.
@@
[…]
@@ Appendix B Proof of Lemma 1
$={\left({\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}}\right)^{2}}\left[{\frac{2}{{r_{i}^{2}}}+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{x_{l}^{2}+y_{l}^{2}+z_{l}^{2}-{{\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right)}^{2}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
Let ${C_{\max}}\triangleq\mathop{\max}\limits_{{{\bf{q}}_{l}}}x_{l}^{2}+y_{l}^{2}+z_{l}^{2}$, it follows that
@@
$\ \ \ \ \ \left[{\frac{2}{{r_{i}^{2}}}+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
With ${\nabla^{2}}{g_{k,i,l}}\preceq{\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{2}}{\bf{I}}$, and by choosing ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{\frac{2}{{r_{i}^{2}}}+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, we have ${\nabla^{2}}{g_{k,i,l}}\preceq{\delta_{k,i,l}}{\bf{I}}$. Thus, the lower and upper bounds of $g_{k,i,l}$}} in (32) and (33) can be obtained with the Taylor’s theorem [38]. The proof of Lemma 1 is thus completed.